Activity at the Soufriere Hills Volcano remained at a relatively low level during this reporting period, very similar to the level in the previous 24 hours. The activity remains concentrated on the north and north-western flanks of the dome and is dominated by small to moderate-sized rockfalls. There was a resumption of the hybrid seismicity that has been associated with dome growth in the past.
A total of 39 rockfall events were recorded during the reporting period, a small decrease compared to yesterday. The largest of these rockfalls occurred at 16:23, 19:28 and 23:21 on 12 June and 09:05 on 13 June. There were 2 long period earthquakes and one VT earthquake, located at a depth of about 1 km beneath the crater. Intermittent low-amplitude broadband tremor was recorded on the Gages seismometer from about 04:30 to 10:00 on 13 June. This is probably due to increased steam emission from the volcano after overnight rain.
The Gages and Chances Peak seismic stations started to record a few small repetitive hybrid events very early in the morning of 13 June. These slowly increased in number and were occuring at a rate of about one every two minutes by the end of the reporting period.
Viewing conditions were generally poor today due to cloud cover. It was however possible to see the lower flanks of the dome on occasions. There was a moderate-sized rockfall on the eastern flank of the dome at 10:06, which generated a small ash cloud. Few other rockfalls were seen. Material continues to build up above Gages Wall. So far, only a small amount has come over the wall and this has started to fill in a small area of the Upper Gages Valley about 300 ft downslope. Material continues to be deposited behind Farrell's Wall; the moat there is now about 75 feet deep. The small spine on the northeasten flank of the dome has collapsed.
EDM measurements were made today on the eastern triangle. The Long Ground to Castle Peak line shows a shortening of 9 mm since the last measurement on 11 June and the Whites to Castle Peak line shows a shortening of 6 mm for the same period. These results continue the recent trend of movement observed for Castle Peak.
The GPS survey has been restarted after a succesfull test of the replacement equipment. No results are yet available.
No COSPEC measurements were made today. Yesterday's measurements show an average SO2 flux of 168 tonnes per day, which is not significantly different from recent results.
Laurance Donnelly of the BGS left Montserrat on Tuesday after spending two months working on the GPS campaign. He has been replaced by Mark Davies, of the Open University, UK, who arrived yesterday.
Scientists at the Montserrat Volcano Observatory remain highly concerned about the current state of the volcano and the dangers to people and property on both the eastern and upper western flanks of the volcano. Visits to the evacuated zone should be kept to an absolute minimum. The Tar River, Long Ground and Whites areas to the east and upper Fort Ghaut, Gages Village and Upper Amersham areas to the west are all now extremely dangerous. The road between Lee's Estate and Ryners Village remains closed. People should not enter these areas under any circumstances. If they do, they put themselves and others at direct risk of very serious injury or horrible death.